OpenStudy (anonymous):

x.5 - 1=9 what does x=?

OpenStudy (anonymous):

is that x * 5? or x + 5 or x - 5 or x/5?

OpenStudy (anonymous):

x/5

OpenStudy (anonymous):

x/5 - 1 = 9 add one to each side x/5 = 10 then multiply each side by 5

OpenStudy (anonymous):

so it would be 50/5?

OpenStudy (anonymous):

\[\frac{ x }{ 5 } * \frac{ 5 }{ 1 } = 10 * 5\]

OpenStudy (anonymous):

cross cancel on the left side to get what?

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):